### v1

@@ 2 The Ht-product of third-order tensors over $\mathbb{H}_{c}$
$=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})(\textbf{Vec}({\mathcal{A}}_{1,i}^{\ast})+j\textbf{Vec}({\mathcal{A}}_{j,k}^{\ast}))$
$=\textbf{Vec}(\widehat{{\mathcal{A}}_{1,i}^{\ast}})+j\textbf{Vec}(\widehat{{\mathcal{A}}_{j,k}^{\ast}}),$
@@
$\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,i}^{\ast}})+j\textbf{diag}(\widehat{{\mathcal{A}}_{j,k}^{\ast}}).$ (5)
Similarly, we have $\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,i}})+j\textbf{diag}(\widehat{{\mathcal{A}}_{j,k}})\in\mathbb{H}_{c}^{n_{3}n_{1}\times n_{3}n_{2}}.$ Its conjugate transpose is $(\textbf{diag}(\hat{\mathcal{A}}))^{\ast}=(\textbf{diag}(\widehat{{\mathcal{A}}_{1,i}}))^{\ast}+j(\textbf{diag}(\widehat{{\mathcal{A}}_{j,k}}))^{\ast}.$ Combing with equality (5) and Lemma 2.12, we have $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$blacksquare
@@ 3 The Ht-SVD of third-order tensors over $\mathbb{H}_{c}$
The SVD of a matrix over $\mathbb{H}_{c}$ has been discussed in some papers (see, e.g., [13]). Let $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n}$. If
@@
where $U=U_{1}e_{1}+U_{2}e_{2}$ and $V=V_{1}e_{1}+V_{2}e_{2}$ are unitary matrices, $\Sigma=\Sigma_{1}e_{1}+\Sigma_{2}e_{2}$ is diagonal but not a real matrix in general.
Recall that a tensor is called $f$-diagonal if each frontal slice is diagonal ([6]). Next, we propose the Ht-SVD of a reduced biquaternion tensor.
@@ 4 The MP inverses of reduced biquaternion tensors
We refer the readers to [1] for Moore-Penrose (MP) inverses. In this section, as an application of Ht-SVD of a RB tensor, we define the MP inverse for a RB tensor and find a formula for the MP inverse by using Theorem 3.1. Moreover, we explore some theoretical properties of the MP inverse of RB tensors.
@@
Now, we introduce the definition of the MP inverse of a third-order Segre quaternion tensor.
@@ 5 Reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$
We first discuss the consistency of RB tensor equation
@@
by using the MP inverse demonstrated in Section 4. The solvability conditions are found and a general expression of the solutions is provided when the solutions exist. Furthermore, we also derive solvability conditions and the general solutions involving Hermicity.
@@
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$ Then equation (9) has solutions if and only if $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$.
Furthermore, the general solutions to equation (9) can be written as
@@
Moreover, each solution $\mathcal{X}$ of equation (9) has the form of the equation (10) since $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{X}.$blacksquare
@@
Let $\mathcal{A},\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then equation (9) has a Hermitian solution if and only if $\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}.$
Moreover, if (9) is solvable, then the general Hermitian solutions to the Segre quaternion tensor equation (9) can be written as
@@
is a Hermitian solution to Segre quaternion tensor equation (9).
Next we show that each solution of Segre quaternion tensor equation (9) should be in the form of (11). Assume $\mathcal{X}_{0}$ is an arbitrary Hermitian solution to tensor equation (9). Setting $\mathcal{U}=\mathcal{X}_{0}$ yields
@@
which implies that each Hermitian solution to Segre quaternion tensor equation (9) can be represented as (11). Thus, (11) is the general expression of Hermitian solutions to equation (9). $\blacksquare$
Next, we deal with the least-squares problem of $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$. It is well known that Segre quaternion matrices have the following properties. The equivalent complex matrix of reduced biquaternion matrix $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n}$ ([13]) is
$\widetilde{A}_{(c)}=\begin{bmatrix}A_{(c),1}&0\\ 0&A_{(c),2}\\ \end{bmatrix}.$ (12)
Given $A,B\in\mathbb{H}_{c}^{m\times n}$ and $C\in\mathbb{H}_{c}^{n\times l}$, we have $(\widetilde{A+B})_{(c)}=\widetilde{A}_{(c)}+\widetilde{B}_{(c)}$ and $(\widetilde{AC})_{(c)}=\widetilde{A}_{(c)}\widetilde{C}_{(c)}$. For $A=(a_{i_{1}i_{2}})\in\mathbb{H}_{c}^{m\times n}$, we define the Frobenius norm of $A$ as $\|A\|_{F}=\sqrt{\sum_{i_{1}=1}^{m}\sum_{i_{2}=1}^{n}|a_{i_{1}i_{2}}|^{2}}$, and $\infty$-norm as $\|A\|_{\infty}=\mathop{\max}_{1\leq i\leq m}\sum_{j=1}^{n}|a_{ij}|$. Moreover, it can be verified that
$\|A\|_{F}=\frac{1}{\sqrt{2}}\sqrt{\|A_{(c),1}\|^{2}+\|A_{(c),2}\|^{2}}=\frac{1}{\sqrt{2}}\|\widetilde{A}_{(c)}\|_{F}.$ (13)
$\|A\|_{\infty}=\mathop{\max}_{1\leq i\leq n_{1}n_{3}}\sum_{j=1}^{{n_{2}}}(|c_{ij}|+|d_{ij}|),$
where $A_{(c),1}=(c_{ij})_{m\times n},$ $A_{(c),2}=(d_{ij})_{m\times n}.$
For a tensor $\mathcal{A}=\mathcal{A}_{1,i}+j\mathcal{A}_{j,k}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,i},\mathcal{A}_{j,k}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$, we define its Frobenius norm as
$\|\mathcal{A}\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A})\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A}_{1,i})\|_{F}^{2}+\|\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}.$ (14)
$\|\mathcal{A}\|_{\infty}=\mathop{\max}_{1\leq i\leq n_{1}n_{3}}\sum_{j=1}^{{n_{2}}}(|c_{ij}|+|d_{ij}|),$ (15)
where $\textbf{Vec}(\mathcal{A}_{1,i})=(c_{ij})_{n_{1}n_{3}\times{n_{2}}},$ $\textbf{Vec}(\mathcal{A}_{j,k})=(d_{ij})_{n_{1}n_{3}\times{n_{2}}}.$
@@
Proof.  By applying identities (3), (14) and $F_{n_{3}}j=jF_{n_{3}}$ as well as the invariant property of Frobenius norm under a unitary transformation for complex matrices, we have
@@
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,i})+\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})j\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}$
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,i})\|_{F}^{2}+\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}$
$=n_{3}\|\textbf{Vec}(\mathcal{A}_{1,i})\|_{F}^{2}+n_{3}\|\textbf{Vec}(\mathcal{A}_{j,k})\|_{F}^{2}$
@@
Lemma 5.4
([1]) The solutions of least-squares problem of complex equation $AX=B$ are
$X=A^{\dagger}B+(I-A^{\dagger}A)Y,$
with $Y$ be any matrix with appropriate size. Furthermore, $A^{\dagger}B$ is the minimum norm least-squares solution.
Lemma 5.5
Suppose $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n},\ B=B_{(c),1}e_{1}+B_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times l},$ Denote $\dot{B}:=\begin{bmatrix}B_{(c),1}\\ B_{(c),2}\\ \end{bmatrix}.$ Then reduced biquaternion least-squares problem $AX=B$ has a solution $X=X_{(c),1}e_{1}+X_{(c),2}e_{2}\in\mathbb{H}_{c}^{n\times l}$ if and only if the complex least-squares problem $\widetilde{A}_{(c)}\dot{X}=\dot{B}$ has a solution $\dot{X}=\begin{bmatrix}X_{(c),1}\\ X_{(c),2}\\ \end{bmatrix}$, where $\widetilde{A}_{(c)}$ is given by (12).
In addition, if $\dot{X}$ is the minimum norm solution to the complex least-squares problem $\widetilde{A}_{(c)}\dot{X}=\dot{B}$, then $X=\begin{bmatrix}I_{n},&0\\ \end{bmatrix}\dot{X}e_{1}+\begin{bmatrix}0,&I_{n}\\ \end{bmatrix}\dot{X}e_{2}$ is the minimum norm solution to the least-squares problem $AX=B$ over $\mathbb{H}_{c}$.
Proof.  By equation (13), we have
$\|AX-B\|_{F}=\frac{1}{\sqrt{2}}\|\widetilde{A}_{(c)}\widetilde{X}_{(c)}-\widetilde{B}_{(c)}\|_{F},$
where $\widetilde{A}_{(c)},\widetilde{B}_{(c)}$ and $\widetilde{X}_{(c)}$ are defined by (12). Then using the fact
$\|\widetilde{A}_{(c)}\widetilde{X}_{(c)}-\widetilde{B}_{(c)}\|_{F}=\|\widetilde{A}_{(c)}\begin{bmatrix}X_{(c),1}\\ X_{(c),2}\\ \end{bmatrix}-\begin{bmatrix}B_{(c),1}\\ B_{(c),2}\\ \end{bmatrix}\|_{F},$
we can see that $\|AX-B\|_{F}$ and $\|\widetilde{A}_{(c)}\dot{X}-\dot{B}\|_{F}$ arrive the minima at the same time. With Lemma 5.4,
$\dot{X}=(\widetilde{A}_{(c)})^{\dagger}\dot{B}+(I-(\widetilde{A}_{(c)})^{\dagger}\widetilde{A}_{(c)})Y$
is the least-squares solution to $\widetilde{A}_{(c)}\dot{X}=\dot{B}$, in which $Y$ is arbitrary, and $X=\begin{bmatrix}I_{n},&0\\ \end{bmatrix}\dot{X}e_{1}+\begin{bmatrix}0,&I_{n}\\ \end{bmatrix}\dot{X}e_{2}$ are the least-squares solutions to $AX=B.$blacksquare
@@
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}$ with their DFTs $\hat{\mathcal{A}}$ and $-^\hat{\mathcal{B}}$. Denote $-^\textbf{A}\doteq\textbf{diag}(\hat{\mathcal{A}}),\textbf{B}\doteq\textbf{Vec}(\hat{\mathcal{B}}).$ Then $\mathcal{X}\in\mathbb{H}^{n_{2}\times n_{4}\times n_{3}}$ is a solution to the tensor least-squares problem $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$ if and only if $\textbf{X}=\textbf{Vec}(\hat{\mathcal{X}})$ is a solution to the matrix least-squares problem $\textbf{A}\textbf{X}=\textbf{B}.$
Moreover, if X is the minimum norm least-squares solution to $\textbf{A}\textbf{X}=\textbf{B}$, then
$=\textbf{Fold}\left(\frac{1}{\sqrt{n_{3}}}(F_{n_{3}}^{\ast}\otimes I_{n_{2}})\textbf{X}\right)$ X )
is the minimum norm least-squares solution to $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}.$
Proof.  By Theorem 2.3 and Theorem 5.3, we have

### v2

@@ 2 The Ht-product of third-order tensors over $\mathbb{H}_{c}$
$=\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{2}})(\textbf{Vec}({\mathcal{A}}_{1,\mathbf{i}}^{\ast})+\mathbf{j}\textbf{Vec}({\mathcal{A}}_{\mathbf{j},\mathbf{k}}^{\ast}))$
$=\textbf{Vec}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}^{\ast}})+\mathbf{j}\textbf{Vec}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}^{\ast}}),$
@@
$\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}^{\ast}})+\mathbf{j}\textbf{diag}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}^{\ast}}).$ (5)
Similarly, we have $\textbf{diag}(\widehat{\mathcal{A}})=\textbf{diag}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}})+\mathbf{j}\textbf{diag}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}})\in\mathbb{H}_{c}^{n_{3}n_{1}\times n_{3}n_{2}}.$ Its conjugate transpose is $(\textbf{diag}(\hat{\mathcal{A}}))^{\ast}=(\textbf{diag}(\widehat{{\mathcal{A}}_{1,\mathbf{i}}}))^{\ast}+\mathbf{j}(\textbf{diag}(\widehat{{\mathcal{A}}_{\mathbf{j},\mathbf{k}}}))^{\ast}.$ Combing with the identity (5) and Lemma 2.12, we have $\textbf{diag}(\widehat{{\mathcal{A}}^{\ast}})=(\textbf{diag}(\widehat{\mathcal{A}}))^{\ast}.$blacksquare
@@ 3 The Ht-SVD of third-order tensors over $\mathbb{H}_{c}$
The SVD of a matrix over $\mathbb{H}_{c}$ has been discussed in some papers (see, e.g., [11]). Let $A=A_{(c),1}e_{1}+A_{(c),2}e_{2}\in\mathbb{H}_{c}^{m\times n}$. If $A_{(c),1}$ and $A_{(c),2}$ have the following SVDs:
@@
where $U=U_{1}e_{1}+U_{2}e_{2}$ and $V=V_{1}e_{1}+V_{2}e_{2}$ are unitary matrices, $\Sigma=\Sigma_{1}e_{1}+\Sigma_{2}e_{2}$ is diagonal but not a real matrix in general. We define the rank of a matrix $A\in\mathbb{H}_{c}^{m\times n}$ is the number of its non-zero singular values.
Recall that a third-order tensor is called $f$-diagonal if each frontal slice is diagonal ([5]). Next, we propose the Ht-SVD of a reduced biquaternion tensor.
@@
Based on Theorem 3.1, we give the reduced biquaternion tensor SVD decomposition (Ht-SVD) algorithm as follows. In MATLAB command, for the complex tensor $\mathcal{C}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$, the DFT of $\mathcal{C}$ and its inverse function are denoted as:
$\hat{\mathcal{C}}=\text{fft}(\mathcal{C},\left[\,\right],3),\ \ \mathcal{C}=\text{ifft}(\hat{\mathcal{C}},\left[\,\right],3).$
Algorithm 1 (SVD-RBT) SVD for a Reduced Biquaternion Tensor
1:  Input: $\mathcal{A}_{(c),1},\mathcal{A}_{(c),2}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$ from given $\mathcal{A}=\mathcal{A}_{(c),1}e_{1}+\mathcal{A}_{(c),2}e_{2}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$
2:  Output: $\mathcal{U}\in\mathbb{H}_{c}^{n_{1}\times n_{1}\times n_{3}},\mathcal{S}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{V}\in\mathbb{H}_{c}^{n_{2}\times n_{2}\times n_{3}},$ such that
$\mathcal{A}=\mathcal{U}\ast_{Ht}\mathcal{S}\ast_{Ht}\mathcal{V}^{\ast}$
3:  $\hat{\mathcal{A}}_{(c),1}$=fft$(\mathcal{A}_{(c),1},[],3)$;$\hat{\mathcal{A}}_{(c),2}$=fft$(\mathcal{A}_{(c),2},[],3)$
4:  Calculate each frontal slice of $\hat{\mathcal{U}},\hat{\mathcal{S}}$ and $\hat{\mathcal{V}}$ from $\hat{\mathcal{A}}$:for $i=1,...n_{3},$ do$[\hat{U}_{1}^{(i)},\hat{S}_{1}^{(i)},\hat{V}_{1}^{(i)}]=svd(\hat{\mathcal{A}}_{(c),1}^{(i)})$, $[\hat{U}_{2}^{(i)},\hat{S}_{2}^{(i)},\hat{V}_{2}^{(i)}]=svd(\hat{\mathcal{A}}_{(c),2}^{(i)})$end for
[…]
6:  $\mathcal{U}=\mathcal{U}_{1}e_{1}+\mathcal{U}_{2}e_{2};$ $\mathcal{S}=\mathcal{S}_{1}e_{1}+\mathcal{S}_{2}e_{2};$ $\mathcal{V}=\mathcal{V}_{1}e_{1}+\mathcal{V}_{2}e_{2};$
7:  return $\mathcal{U}$, $\mathcal{S}$, $\mathcal{V}$
Complexity Analysis of Algorithm 1: For a given reduced biquaternion tensor $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$, the computational complexity of each iteration of Algorithm 1 (SVD-RBT) is primarily driven by two major components: the fast fourier transform (FFT) and its inverse, and the SVD of the tensor’s frontal slices. The computational complexity of FFT and inverse FFT operations is $O(n_{1}n_{2}n_{3}\log n_{3})$. The SVD process, applied separately to each frontal slice, has a complexity of $O(n_{3}\min(n_{1}^{2}n_{2},n_{1}n_{2}^{2}))$. Additionally, the tensor reconstruction step, which involves simple element-wise addition, has a complexity of $O(n_{1}n_{2}n_{3})$. Thus, the total computational complexity for each iteration of the algorithm is $O(n_{1}n_{2}n_{3}\log n_{3}+n_{3}\min(n_{1}^{2}n_{2},n_{1}n_{2}^{2}))$.
@@ 4 The MP inverses of reduced biquaternion tensors
We refer the readers to [1] for Moore-Penrose (MP) inverses. In this section, as an application of Ht-SVD of a RB tensor, we define the MP inverse for a RB tensor and find a formula for the MP inverse by using Theorem 3.1. Moreover, we explore some theoretical properties of the MP inverses of RB tensors.
@@
Now, we introduce the definition of the MP inverse of a third-order reduced biquaternion tensor.
@@ 5 Reduced biquaternion tensor equation $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$
We mainly discuss the consistency of RB tensor equation
@@
by using the MP inverse demonstrated in Section 4. The solvability conditions are found and a general expression of the solutions is provided when the solutions exist. Except that, we derive the least-squares solutions/the minimal norm least-squares solution to (9). Furthermore, we also derive solvability conditions and the general solutions involving Hermicity. One of the applications of solving this tensor equations is color video deblurring given in Section 6.
@@
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$ Then equation (9) has solutions if and only if $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}$. Furthermore, the general solutions to equation (9) can be written as
@@
Next we show each solution $\mathcal{X}$ of equation (9) has the form of $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+\mathcal{L}_{\mathcal{A}}\ast_{Ht}\mathcal{X}$. Note that $\mathcal{A}\ast_{Ht}\mathcal{X}=\mathcal{B}$, thus $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A}\ast_{Ht}\mathcal{X}+\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-{\mathcal{A}}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{X},$ as required. Based on the above, we can conclude that (10) is the general expression of the solutions. $\blacksquare$
@@
Let $\mathcal{A},\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$. Then equation (9) has a Hermitian solution if and only if $\mathcal{A}\ast_{Ht}\mathcal{B}^{\ast}=\mathcal{B}\ast_{Ht}\mathcal{A}^{\ast}$ and $\mathcal{R}_{\mathcal{A}}\ast_{Ht}\mathcal{B}=\mathcal{O}.$ Moreover, if (9) is solvable, then the general Hermitian solutions to the reduced biquaternion tensor equation (9) can be written as
@@
is a Hermitian solution to reduced biquaternion tensor equation (9).
Next we show that each solution of reduced biquaternion tensor equation (9) should be in the form of (11). Assume $\mathcal{X}_{0}$ is an arbitrary Hermitian solution to tensor equation (9). Setting $\mathcal{U}=\mathcal{X}_{0}$ yields
@@
which implies that each Hermitian solution to reduced biquaternion tensor equation (9) can be represented as (11). Thus, (11) is the general expression of Hermitian solutions to equation (9). $\blacksquare$
Next, we deal with the least-squares problem of RB tensor equation (9). For a tensor $\mathcal{A}=\mathcal{A}_{1,\mathbf{i}}+\mathbf{j}\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ with $\mathcal{A}_{1,\mathbf{i}},\mathcal{A}_{\mathbf{j},\mathbf{k}}\in\mathbb{C}^{n_{1}\times n_{2}\times n_{3}}$, we define its Frobenius norm as
$\|\mathcal{A}\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A})\|_{F}^{2}=\|\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})\|_{F}^{2}+\|\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}.$ (12)
where $\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})=(c_{ij})_{n_{1}n_{3}\times{n_{2}}},$ $\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})=(d_{ij})_{n_{1}n_{3}\times{n_{2}}}.$
@@
Proof.  By applying identities (3), (12) and $F_{n_{3}}\mathbf{j}=\mathbf{j}F_{n_{3}}$ as well as the invariant property of Frobenius norm under a unitary transformation for complex matrices, we have
@@
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})+\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\mathbf{j}\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}$
$=\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})\|_{F}^{2}+\|\sqrt{n_{3}}(F_{n_{3}}\otimes I_{n_{1}})\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}$
$=n_{3}\|\textbf{Vec}(\mathcal{A}_{1,\mathbf{i}})\|_{F}^{2}+n_{3}\|\textbf{Vec}(\mathcal{A}_{\mathbf{j},\mathbf{k}})\|_{F}^{2}$
@@
Let $\mathcal{A}\in\mathbb{H}_{c}^{n_{1}\times n_{2}\times n_{3}}$ and $\mathcal{B}\in\mathbb{H}_{c}^{n_{1}\times n_{4}\times n_{3}}.$ Then the least-squares solutions of tensor equation (9) can be represented as $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}+(\mathcal{I}-\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{A})\ast_{Ht}\mathcal{W}$, where $\mathcal{W}$ is an arbitrary compatible RB tensor, and the minimum-norm least-squares solution of tensor equation (9) is $\mathcal{X}=\mathcal{A}^{\dagger}\ast_{Ht}\mathcal{B}$.
Proof.  According to Theorem 2.3 and Theorem 5.3, we have